# On an Optimal Stopping Problem with a Discontinuous Reward

### v3

@@ 4.3. Surrender and Continuation Premium Representation
In this section and throughout the rest of the paper, derivatives should be understood as distributional.
@@
(24) $e(t,x):=\int_{t}^{T}\Big{(}c(s)g(s)-g_{t}(s)\Big{)}\mathbb{E}\Big{[}e^{-r(s-t)}F_{s}^{t,x}\mathbf{1}_{\{(s,F_{s}^{t,x})\in\mathcal{S}\}}\Big{]}\,\mathrm{d}s,$
where $g_{t}=\frac{\partial g}{\partial t}$.
@@
Corollary 4.2 ensures that $v$ is smooth enough to apply Itô’s formula for generalized derivatives (see Krylov (1980), Theorem 2.10.1) to $\left\{e^{-rs}v(t,F_{s}^{t,x})\right\}_{t\leq s\leq T}$, yielding
@@
for $0\leq s\leq T$. Integrating from $t$ to $T$ on both side and multiplying by $e^{rt}$, we get
@@
Now, ${\mathbb{E}\left[e^{-r(T-t)}v(T,F_{T}^{t,x})\right]=\mathbb{E}\left[e^{-r(T-t)}\max(G,F_{T}^{t,x})\right]=h(t,\,x)}$, and by Corollary 4.1, $\mathcal{L}_{t}v(s,x)+v_{t}(s,x)-rv(s,x)=0$ for all $(s,\,x)\in\mathcal{C}$. Futhermore, for $s<T$, $v(s,\,x)=g(s)x$ when $(s,x)\in\mathcal{S}$. It follows that
[…]
which concludes the proof.
@@
(26) $e(t,x):=\int_{t}^{T}\Big{(}c(s)g(s)-g_{t}(s)\Big{)}\mathbb{E}\Big{[}e^{-r(s-t)}F_{s}^{t,x}\mathbf{1}_{\{F_{s}^{t,x}>b(s)\}}\Big{]}\,\mathrm{d}s.$
The results presented in Proposition 4.4 are generalizations of Theorem 1 and Equation (11) of Bernard et al. (2014b) to time-dependent fee and surrender charge functions.
@@
Proposition 4.5 decomposes the value function in terms of the immediate surrender value plus a term representing the value of holding on to the contract. This term, which we coin the continuation premium, is the sum of the financial guarantee at maturity (or the put option) when the contract is held until $T$ and an integral term equal to the value added by keeping the contract until it is optimal to surrender. The representation of the continuation premium in (28) can be particularly helpful to develop numerical methods for approximating the value of a variable annuity contract, since it uses additional information on the shape of the continuation region.
It is also possible to decompose the value function into the surrender benefit and the continuation premium in a more general setting, where Assumptions 4.1 and 4.2 do not need to hold. However, in this case, the continuation premium is written in terms of $\tau_{t}^{x}$, which can make it more complicated to calculate. This result is presented in Lemma 4.3 below and will be used to prove Proposition 4.5.
@@
The value function (7) can be written as
(30) $v(t,x)=xg(t,x)+f(t,x),$
@@
with ${Y^{t,x}=\{Y_{s}^{t,x}\}_{t\leq s\leq T}}$, with $Y_{s}^{t,x}:=e^{-r(s-t)}g(s,F_{s}^{t,x})F_{s}^{t,x}$ and observe that ${Y_{T}^{t,x}=e^{-r(T-t)}F_{T}^{t,x}}$ since ${g(T,x)=1}$ for all $x\in\mathbb{R}_{+}$. The function $xg(t,x)$ is $C^{1,2}$, so we can apply Itô’s formula to $Y$, which yields
@@
(34) $\begin{split}&\mathbb{E}\left[e^{-r(T-t)}(G-F_{T}^{t,x})_{+}\mathbf{1}_{\{\tau_{t}^{x}\neq T\}}\right]\\ &\qquad+\mathbb{E}\left[\int_{\tau_{t}^{x}}^{T}\Big{(}g_{t}(s)-c(s)g(s)\Big{)}e^{-r(s-t)}F_{s}^{t,x}\mathbf{1}_{\{(s,F_{s})\in\mathcal{C}\}}\,\mathrm{d}s\right]=0,\end{split}$
@@
$=xg(t)+\mathbb{E}\left[e^{-r(T-t)}(G-F_{T}^{t,x})_{+}\right]+\int_{t}^{T}L(s)\mathbb{E}\Big{[}e^{-r(s-t)}F_{s}^{t,x}\mathbf{1}_{\{F_{s}^{t,x}\in\mathcal{C}_{s}\}}\Big{]}\,\mathrm{d}s,$
@@ 4.4. Characterization of the Surrender Region
    If $L(t)<0$ for all $t\in[0,T)$ then the $t$-sections of $\mathcal{S}$, denoted by $\mathcal{S}_{t}$, are of the form
    (39) $\mathcal{S}_{t}=\{x\in\mathbb{R}_{+}:(t,x)\in\mathcal{S}\}=[b(t),\,\infty),$
@@
    If $L(t)\geq 0$ for all $(t)\in[0,T)$, then $\mathcal{S}=\emptyset$ and $T$ is the unique optimal stopping time for (7).
@@
respectively. When $b(t)<\infty$ for all $t\in[0,T)$, the surrender boundary splits $[0,T)\times\mathbb{R}_{+}$ in two regions: the surrender region is at or above the boundary, and the continuation region is below. It follows that the set $\mathcal{S}$ is connected. Henceforth, we say that the surrender region is of “threshold type” if for any $t\in[0,T)$, there exists a $b(t)<\infty$ such that $\mathcal{S}_{t}=[b(t),\infty)$. Such a geometry for the surrender region can be explained, as in Milevsky and Salisbury (2001), by the fact that when the account value is low, it is financially advantageous for the policyholder to hold on to the contract since there is a significant chance that the guarantee will be triggered at maturity. Since this guarantee is financed by the policyholder via the continuous fee, which reduces the net return on the account, there is a point above which it is no longer profitable to hold the contract and continue paying the fee; this threshold is the optimal surrender boundary.
@@
Proposition 4.6 characterizes the non-emptiness of the surrender region. The first part involves $\text{relint}(\mathcal{S}_{t})$, the relative interior of a $t$-section $\mathcal{S}_{t}$, which is defined as its interior within the affine hull of $\mathcal{S}_{t}$. If it is not empty, the relative interior of $\mathcal{S}_{t}$ forms a vertical line in the plane. If $\text{relint}(\mathcal{S}_{t})\in\text{int}(\mathcal{S})$, it contains some finite $x$ for which $(t,x)\in\text{int}(\mathcal{S})$. The second part of Proposition 4.6 (ii) can be seen as a local version of the first part.
@@
      $\mathcal{S}\neq\emptyset$ and $\mathcal{C}\neq\emptyset$, and
@@
      $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})\neq\emptyset$ for any $[t_{1},t_{2})\subset[0,T)$ such that $t_{1}<t_{2}$.
@@
    From Theorem 4.2, Remark 4.9 and Corollary 4.2, we have that $v\in C^{0,1}([0,T]\times\mathbb{R}_{+}\setminus\{T\}\times\{G\})\cap C^{1,2}([0,T]\times\mathbb{R}_{+}\setminus\partial\mathcal{C})$, so that $v\in C^{1,2}(\text{int}(\mathcal{S}))$. Observe that for $x\in\text{relint}(\mathcal{S}_{t})$, $v(t,x)=xg(t)$, so that $xL(t)=\mathcal{L}_{t}v+v_{t}-rv$. We note that the differential operator can be applied since $x\in\int(\mathcal{S})$. Recall from Proposition 4.2 that under Assumption 4.2, $\mathcal{L}_{t}v+v_{t}-rv\leq 0$ for all $(t,x)\in(0,T)\times\mathbb{R}_{+}$. The result follows since $x>0$ in $\text{int}(\mathcal{S})$.
@@
    To show the first statement, we use the continuation premium representation (28), so that
@@
    for all $(t,x)\in[0,T]\times\mathbb{R}_{+}$. We want to show that there exists some $(t_{1},x_{1})\in[0,T)\times\mathbb{R}_{+}$ satisfying $v(t_{1},x_{1})=\varphi(t_{1},x_{1})$ and some $(t_{2},x_{2})\in[0,T)\times\mathbb{R}_{+}$ satisfying ${v(t_{2},x_{2})>\varphi(t_{2},x_{2})}$. We proceed by contradiction.
    First, suppose that $L(s)<0$ for all $s\in[0,T)$ and $\mathcal{S}=\emptyset$. Fix $t\in[0,T)$. Since $\mathcal{S}=\emptyset$, we have $\mathcal{C}=[0,T)\times\mathbb{R}_{+}$, so that (41) becomes
    $v(t,x)-\varphi(t,x)$
    $\quad=\mathbb{E}\left[e^{-r(T-t)}(G-F_{T}^{t,x})_{+}\right]+\int_{t}^{T}\left(g_{t}(s)-c(s)g(s)\right)xe^{-\int_{t}^{s}c(u)\,\mathrm{d}u}\,\mathrm{d}s$
    (42) $\quad=\mathbb{E}\left[e^{-r(T-t)}(G-F_{T}^{t,x})_{+}\right]+x\left(e^{-\int_{t}^{T}c(s)\,\mathrm{d}s}-g(t)\right).$
    Now note that $\mathcal{S}_{t}=\emptyset$ implies $v(t,x)-\varphi(t,x)>0$ for all $x\in\mathbb{R}_{+}$. Hence, we deduce that for all $x\in\mathbb{R}_{+}$,
    (43) $\mathbb{E}\left[e^{-r(T-t)}(G-F_{T}^{t,x})_{+}\right]>x\left(g(t)-e^{-\int_{t}^{T}c(s)\,\mathrm{d}s}\right).$
    Observe that $L(s)=g_{t}(s)-c(s)g(s)<0\Leftrightarrow\frac{\,\mathrm{d}\ln g(s)}{\,\mathrm{d}s}<c(s)$ for all $s\in[0,T)$. By integrating from $t$ to $T$ on both sides of the inequality and using the fact that $g(T)=1$, we find that
    (44) $g(t)-e^{-\int_{t}^{T}c(s)\,\mathrm{d}s}>0,$
    for all $t\in[0,T)$.
    The expectation on the left-hand side of (43) is a continuous and monotonically decreasing function of $x$, is equal to $Ge^{-r(T-t)}$ when $x=0$ and approaches 00 as $x\rightarrow\infty$. The function on the right-hand side is continuous (under Assumption 4.1), strictly increasing in $x$ and ranges from 0 to $\infty$. Then, there must exist $y\in\mathbb{R}_{+}$ such that $\mathbb{E}\left[e^{-r(T-t)}(G-F_{T}^{t,x})_{+}\right]\leq x\left(e^{-\int_{t}^{T}c(s)\,\mathrm{d}s}-g(t)\right)$ for all $x\geq y$, contradicting (43). We conclude that ${\mathcal{S}_{t}\neq\emptyset}$.
    For the second part of the proof, we proceed similarly. Suppose that $L(s)<0$ for all $s\in[0,T)$ and $\mathcal{C}=\emptyset$. Fix $t\in[0,T)$. Since $\mathcal{C}=\emptyset$, (41) becomes
@@
    Now note that $\mathcal{C}_{t}=\emptyset$ implies $v(t,x)-\varphi(t,x)=0$ for all $x\in\mathbb{R}_{+}$. However, we know that for all $x\in\mathbb{R}_{+}$, $\mathbb{E}\left[e^{-r(T-t)}(G-F_{T}^{t,x})_{+}\right]>0$, so that $\mathcal{C}_{t}\neq\emptyset$, leading to a contradiction. Thus, we conclude that $\mathcal{C}_{t}\neq\emptyset$.
    To show that $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})\neq\emptyset$ for any $[t_{1},t_{2})\subset[0,T)$, we proceed by contradiction. Fix $0\leq t_{1}<t_{2}<T$. Suppose that $L(s)<0$ for all $s\in[0,T)$ and $\cup_{t_{1}\leq t<t_{2}}\mathcal{S}_{t}=\emptyset$ This implies $\mathcal{C}_{t}=\mathbb{R}_{+}$ for all $t\in[t_{1},t_{2})$, so that $v(t,x)-\varphi(t,x)>0$ for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$. Fix $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$. Using the continuation premium representation (28) with $\mathcal{C}_{t}=\mathbb{R}_{+}$ for all $t\in[t_{1},t_{2})$, we have
    $\begin{split}v(t,x)-\varphi(t,x)&=\mathbb{E}\left[e^{-r(T-t)}\left(G-F_{T}^{t,x}\right)_{+}\right]\\ &\quad+\int_{t}^{T}L(s)\mathbb{E}\left[e^{-r(s-t)}F_{s}^{t,x}\right]\,\mathrm{d}s>0,\end{split}$
    so that,
    $\mathbb{E}\left[e^{-r(T-t)}\left(G-F_{T}^{t,x}\right)_{+}\right]$ $>-\int_{t}^{T}L(s)\mathbb{E}\left[e^{-r(s-t)}F_{s}^{t,x}\right]\,\mathrm{d}s$
    (45) $=-x\sup_{t\leq r<t_{2}}L(r)\int_{t}^{t_{2}}e^{-\int_{t}^{s}c(u)\,\mathrm{d}u}\,\mathrm{d}s$
    Note that the function on the right-hand side is positive, equal to 0 when $x=0$, strictly increasing in $x$ (since $L(t)<0$ by assumption) and approaches infinity as $x\rightarrow\infty$. Then, using arguments similar to those of the proof of Proposition 4.6 (ii), there must exist ${y\in\mathbb{R}_{+}}$ such that for all ${x\geq y}$, ${\mathbb{E}\left[e^{-r(T-t)}\left(G-F_{T}^{t,x}\right)_{+}\right]\leq-x\sup_{t\leq s<t_{2}}L(s)\int_{t}^{t_{2}}e^{-\int_{t}^{s}c(u)\,\mathrm{d}u}\,\mathrm{d}s}$, contradicting the strict inequality in (45) which must hold for all ${x\in\mathbb{R}_{+}}$. We conclude that or $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})\neq\emptyset$.
@@
    For each $t\in[0,T)$, if $L(t)>0$, then $\mathcal{S}_{t}=\emptyset$ or $\mathcal{S}_{t}\in\partial\mathcal{S}$.
@@
    For any $[t_{1},t_{2})\subset[0,T)$ such that $t_{1}<t_{2}$, if $L(t)>0$ for all $(t)\in[t_{1},t_{2})$, then $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})=\emptyset$.
@@
    If $\mathcal{S}=\emptyset$, then $g(t)-e^{-\int_{t}^{T}c(s)\,\mathrm{d}s}\leq 0$ for all $t\in[0,T)$.
@@
Some of the assumptions of Proposition 4.7 can be relaxed. The proof of part (i) only requires Assumption 4.2. Furthermore, Proposition 4.7(ii) holds in a general setting, without Assumptions 4.1 and 4.2. In this general setting, it holds that for any $[t_{1},t_{2})\subset[0,T)$ such that $t_{1}<t_{2}$, if $L(t,x)>0$ for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$, then $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})=\emptyset$.
@@
    Observe that $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})=\emptyset$ implies ${\mathcal{C}\cap([t_{1},t_{2})\times\mathbb{R}_{+})=[t_{1},t_{2})\times\mathbb{R}_{+}}$. Thus, we need to show that $v(t,x)>\varphi(t,x)$ for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$. Fix $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$ and recall from the proof of Lemma 4.3 that for any $0<t<s\leq T$
@@
    $\mathbb{E}\left[e^{-r(t_{2}-t)}\varphi(t_{2},F_{t_{2}}^{t,x})\right]$ $=xg(t,x)+\int_{t}^{t_{2}}\mathbb{E}[L(s,F_{s}^{t,x})e^{-r(s-t)}F_{s}^{t,x}]\,\mathrm{d}s$
@@
    $>xg(t,x)=\varphi(t,x),$
    where the last inequality follows since $L(t,x)>0$ for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$. It follows that
@@
    Since $\mathcal{S}=\emptyset$, $\mathcal{C}=[0,T)\times\mathbb{R}_{+}$, $v(t,x)-\varphi(t,x)>0$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$. Then, for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, $v(t,x)-\varphi(t,x)>0$ implies
@@
    As observed in the proof of Proposition 4.6 (ii), the expectation on the left-hand side is a continuous and monotonically decreasing function of $x$, is equal to $Ge^{-r(T-t)}$ when $x=0$ and decreases to 00 as $x\rightarrow\infty$. Therefore, in order for the inequality to hold for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, it must be true that
@@
The first statements of Propositions 4.6 and 4.7 can apply to surrender charge functions that depend on both time and the value of the underlying process. Indeed, the proof relies on the results of Proposition 4.2 which only requires the fee function, and not the surrender charge, to depend exclusively on time, see Remark 4.10 for details.

### v4

@@ 4.3. Surrender and Continuation Premium Representation
In this section and throughout the rest of the paper, derivatives are always considered in the weak sense.
@@
(25) $e(t,x):=\int_{t}^{T}\Big(c(s)g(s)-g_{t}(s)\Big)\mathbb{E}\Big[e^{-r(s-t)}F_{s}^{t,x}\mathbf{1}_{\{(s,F_{s}^{t,x})\in\mathcal{S}\}}\Big]\,\mathrm{d}s.$
@@
Theorem 4.2 ensures that $v$ is smooth enough to apply Itô’s formula for generalized derivatives (see Krylov (1980), Theorem 2.10.1) to $\left\{e^{-rs}v(t,F_{s}^{t,x})\right\}_{t\leq s\leq T}$, yielding
@@
for $0\leq s\leq T$. Integrating from $t$ to $T$ on both sides and multiplying by $e^{rt}$, we get
@@
Now, $\mathbb{E}\left[e^{-r(T-t)}v(T,F_{T}^{t,x})\right]=\mathbb{E}\left[e^{-r(T-t)}\max(G,F_{T}^{t,x})\right]=h(t,\,x)$}, and by Corollary 4.1, $\mathcal{L}_{t}v(s,x)+v_{t}(s,x)-rv(s,x)=0$ for all $(s,\,x)\in\mathcal{C}$. Furthermore, for $s<T$, v(s,x)=g(s)xv(s,\,x)=g(s)x when $(s,x)\in\mathcal{S}$. It follows that
$\mathcal{L}v(s,F_{s}^{t,x})+v_{t}(s,\,F_{s}^{t,x})-rv(s,\,F_{s}^{t,x})=F_{s}^{t,x}\left[g_{t}(s)-c(s)g(s)\right]\mathbf{1}_{\{(s,F_{s}^{t,x})\in\mathcal{S}\}},\quad\textrm{a.s.,}$
which concludes the proof. ∎
@@
(27) $e(t,x):=\int_{t}^{T}\Big(c(s)g(s)-g_{t}(s)\Big)\mathbb{E}\Big[e^{-r(s-t)}F_{s}^{t,x}\mathbf{1}_{\{F_{s}^{t,x}\geq b(s)\}}\Big]\,\mathrm{d}s.$
Proposition 4.3 is a generalization of Theorem 1 and Equation (11) of Bernard et al. (2014b) to time-dependent fee and surrender charge functions.
@@
Proposition 4.4 decomposes the value function in terms of the immediate surrender value plus a term representing the value of holding on to the contract. This term, which we coin the continuation premium, is the sum of the financial guarantee at maturity when the contract is held until $T$ and an integral term equal to the value added by keeping the contract until it is optimal to surrender. The representation in (30) can be particularly helpful when developing numerical methods for approximating the value of a variable annuity contract, since it uses additional information on the shape of the continuation region.
It is also possible to decompose the value function into the surrender benefit and the continuation premium in a more general setting, where Assumptions 4.1 and 4.2 do not need to hold. However, in this case, the continuation premium is written in terms of $\tau_{t}^{x}$, which complexifies computations. This result is presented in Lemma 4.3 below and is used in the proof of Proposition 4.4.
@@
The continuation premium in Proposition 4.4 can be written as
@@
with $Y^{t,x}=\{Y_{s}^{t,x}\}_{t\leq s\leq T}$}, with $Y_{s}^{t,x}:=e^{-r(s-t)}g(s,F_{s}^{t,x})F_{s}^{t,x}$ and observe that $Y_{T}^{t,x}=e^{-r(T-t)}F_{T}^{t,x}$} since g(T,x)=1{g(T,x)=1} for all $x\in\mathbb{R}_{+}$. The function $xg(t,x)$ is $C^{1,2}$, so we can apply Itô’s formula to $Y$, which yields
@@
(34) $\begin{split}&\mathbb{E}\left[e^{-r(T-t)}(G-F_{T}^{t,x})_{+}\mathbf{1}_{\{\tau_{t}^{x}\neq T\}}\right]\\ &\qquad+\mathbb{E}\left[\int_{\tau_{t}^{x}}^{T}\Big(g_{t}(s)-c(s)g(s)\Big)e^{-r(s-t)}F_{s}^{t,x}\mathbf{1}_{\{(s,F_{s}^{t,x})\in\mathcal{C}\}}\,\mathrm{d}s\right]=0,\end{split}$
@@
$=xg(t)+\mathbb{E}\left[e^{-r(T-t)}(G-F_{T}^{t,x})_{+}\right]$
$\qquad\qquad+\int_{t}^{T}L(s)\mathbb{E}\Big[e^{-r(s-t)}F_{s}^{t,x}\mathbf{1}_{\{F_{s}^{t,x}\in\mathcal{C}_{s}\}}\Big]\,\mathrm{d}s,$
@@ 4.4. Characterization of the Surrender Region
    If $L(t)<0$ for all $t\in[0,T)$ then the $t$-sections $\mathcal{S}_{t}$ are of the form
    (39) $\mathcal{S}_{t}=[b(t),\,\infty),$
@@
    If $L(t)\geq 0$ for all $t\in[0,T)$, then $\mathcal{S}=\emptyset$ and $T$ is the unique optimal stopping time for (7).
@@
respectively. When $b(t)<\infty$ for all $t\in[0,T)$, the surrender boundary splits $0,T)\times\mathbb{R}_{+}$ in two regions: the surrender region is at or above the boundary, and the continuation region is below. It follows that the set $\mathcal{S}$ is connected. Henceforth, we say that the surrender region is of “threshold type” if for any $t\in[0,T)$, there exists a $b(t)<\infty$ such that $\mathcal{S}_{t}=[b(t),\infty)$. Such a geometry for the surrender region can be explained, as in Milevsky and Salisbury (2001), by the fact that when the account value is low, it is financially advantageous for the policyholder to hold on to the contract since there is a significant chance that the guarantee will be triggered at maturity. Since this guarantee is financed by the policyholder via the continuous fee, which reduces the net return on the account, there is a point above which it is no longer profitable to hold the contract and continue paying the fee; this threshold is the optimal surrender boundary.
@@
Proposition 4.5 characterizes the non-emptiness of the surrender region. The first part involves $\text{relint}(\mathcal{S}_{t})$, the relative interior of a $t$-section $\mathcal{S}_{t}$, which is defined as its interior within the affine hull of $\mathcal{S}_{t}$. If it is not empty, the relative interior of $\mathcal{S}_{t}$ forms a vertical line in the plane. If $\text{relint}(\mathcal{S}_{t})\in\text{int}(\mathcal{S})$, it contains some finite $x$ for which $(t,x)\in\text{int}(\mathcal{S})$. The first part of Proposition 4.5 (ii) can be seen as a local version of the second part.
@@
      $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})\neq\emptyset$ for any $0\leq t_{1}<t_{2}<T$, and
@@
      $\mathcal{S}\neq\emptyset$ and $\mathcal{C}\neq\emptyset$.
@@
    From the definition of the value function on $\mathcal{S}$ and of the surrender charge function $g$, $v\in C^{1,2}(\text{int}(\mathcal{S}))$. Observe that for $x\in\text{relint}(\mathcal{S}_{t})$, $v(t,x)=xg(t)$, so that $xL(t)=\mathcal{L}_{t}v+v_{t}-rv$. Recall from Proposition 4.2 that $\mathcal{L}_{t}v+v_{t}-rv\leq 0$ for all $(t,x)\in(0,T)\times\mathbb{R}_{+}$. The result follows since $x>0$.
@@
    To show that $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})\neq\emptyset$ for any $0\leq t_{1}<t_{2}<T$, we proceed by contradiction. Fix $0\leq t_{1}<t_{2}<T$. Suppose that $L(s)<0$ for all $s\in[0,T)$ and $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})=\emptyset$ . This implies $\mathcal{C}_{t}=\mathbb{R}_{+}$ for all $t\in[t_{1},t_{2})$, so that $v(t,x)-\varphi(t,x)>0$ for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$. Fix $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$. Using the continuation premium representation (29) with $\mathcal{C}_{t}=\mathbb{R}_{+}$ for all $t\in[t_{1},t_{2})$, we have
    $\begin{split}v(t,x)-\varphi(t,x)&=\mathbb{E}\left[e^{-r(T-t)}\left(G-F_{T}^{t,x}\right)_{+}\right]\\ &\quad+\int_{t}^{T}L(s)\mathbb{E}\left[e^{-r(s-t)}F_{s}^{t,x}\mathbf{1}_{\{F_{s}^{t,x}\in\mathcal{C}_{s}\}}\right]\,\mathrm{d}s>0,\end{split}$
    so that,
    $\mathbb{E}\left[e^{-r(T-t)}\left(G-F_{T}^{t,x}\right)_{+}\right]$ $>-\int_{t}^{T}L(s)\mathbb{E}\left[e^{-r(s-t)}F_{s}^{t,x}\mathbf{1}_{\{F_{s}^{t,x}\in\mathcal{C}_{s}\}}\right]\,\mathrm{d}s$
    (41) $\geq-x\sup_{t\leq r<t_{2}}L(r)\int_{t}^{t_{2}}e^{-\int_{t}^{s}c(u)\,\mathrm{d}u}\,\mathrm{d}s.$
    Note that the function on the right-hand side is positive, equal to 0 when $x=0$, strictly increasing in $x$ (since $L(t)<0$ by assumption) and approaches infinity as $x\rightarrow\infty$. Then, since the left-hand side of (41) is bounded by $Ge^{-r(T-t)}$, there must exist $y\in\mathbb{R}_{+}$} such that for all $x\geq y$}, $\mathbb{E}\left[e^{-r(T-t)}\left(G-F_{T}^{t,x}\right)_{+}\right]\leq-x\sup_{t\leq s<t_{2}}L(s)\int_{t}^{t_{2}}e^{-\int_{t}^{s}c(u)\,\mathrm{d}u}\,\mathrm{d}s$}, contradicting the strict inequality in (41) which must hold for all $x\in\mathbb{R}_{+}$}. We conclude that or $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})\neq\emptyset$.
    The first part of the second statement, which states that $\mathcal{S}$ is not empty, is obtained by taking $t_{1}=0$ and letting $t_{2}=\lim_{\epsilon\rightarrow 0^{+}}T-\epsilon$ in the first statement. To show that the continuation region is not empty either, we use the continuation premium representation (29), so that
@@
    for all $(t,x)\in[0,T]\times\mathbb{R}_{+}$. We want to show that there exists some $(t_{1},x_{1})\in[0,T)\times\mathbb{R}_{+}$ satisfying $v(t_{2},x_{2})>\varphi(t_{2},x_{2})$}. We proceed by contradiction.
    Suppose that $L(s)<0$ for all $s\in[0,T)$ and $\mathcal{C}=\emptyset$. Fix $t\in[0,T)$. Since $\mathcal{C}=\emptyset$, (42) becomes
@@
    Now note that $\mathcal{C}_{t}=\emptyset$ implies $v(t,x)=\varphi(t,x)$ for all $x\in\mathbb{R}_{+}$. However, we know that for all $x\in\mathbb{R}_{+}$, $\mathbb{E}\left[e^{-r(T-t)}(G-F_{T}^{t,x})_{+}\right]>0$, leading to a contradiction. Thus, we conclude that $\mathcal{C}_{t}\neq\emptyset$, and thus $\mathcal{C}\neq\emptyset$.
@@
    For each $t\in[0,T)$, if $L(t)>0$, then $\mathcal{S}_{t}=\emptyset$ or $\mathcal{S}_{t}\subseteq\partial\mathcal{S}$.
@@
    For any $t_{1},t_{2})\subset[0,T)$ such that $t_{1}<t_{2}$, if $L(t)>0$ for all $t\in[t_{1},t_{2})$, then $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})=\emptyset$.
@@
    If $\mathcal{S}=\emptyset$, then $g(t)\leq e^{-\int_{t}^{T}c(s)\,\mathrm{d}s}$ for all $t\in[0,T)$.
@@
Some of the assumptions of Proposition 4.6 can be relaxed. The proof of part (i) only requires Assumption 4.2. Furthermore, part (ii) holds in a general setting, without Assumptions 4.1 and 4.2. In this general setting, it holds that for any $t_{1},t_{2})\subset[0,T)$ such that $t_{1}<t_{2}$, if $L(t,x)>0$ for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$, then $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})=\emptyset$.
@@
    Since $\mathcal{S}\cap([t_{1},t_{2})\times\mathbb{R}_{+})=\emptyset$ implies $\mathcal{C}\cap([t_{1},t_{2})\times\mathbb{R}_{+})=[t_{1},t_{2})\times\mathbb{R}_{+}$}, we need to show that $v(t,x)>\varphi(t,x)$ for all $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$. Fix $(t,x)\in[t_{1},t_{2})\times\mathbb{R}_{+}$ and recall from the proof of Lemma 4.3 that for any $0<t<s\leq T$
@@
    $\mathbb{E}\left[e^{-r(t_{2}-t)}\varphi(t_{2},F_{t_{2}}^{t,x})\right]$ $=xg(t)+\int_{t}^{t_{2}}\mathbb{E}[L(s,F_{s}^{t,x})e^{-r(s-t)}F_{s}^{t,x}]\,\mathrm{d}s$
@@
    $>xg(t)=\varphi(t,x),$
    where the last inequality follows since $L(t)>0$ for all $t\in[t_{1},t_{2})$. It follows that
@@
    Since $\mathcal{S}=\emptyset$, $\mathcal{C}=[0,T)\times\mathbb{R}_{+}$, $v(t,x)-\varphi(t,x)>0$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, which implies
@@
    As observed in the proof of Proposition 4.5 (ii), the expectation on the left-hand side is a continuous and monotonically decreasing function of $x$, is equal to $Ge^{-r(T-t)}$ when $x=0$ and decreases to 0 as $x\rightarrow\infty$. Therefore, in order for the inequality to hold for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, it must be true that
@@
The first statements of Propositions 4.5 and 4.6 can apply to surrender charge functions that depend on both time and the value of the underlying process.